Can some please help me solve this system of equations?
5x + 3y = 1
x + 2y = 10.

Answer by dabanfield(803)

Can some please help me solve this system of equations?
1.) 5x + 3y = 1
2.) x + 2y = 10.
Let's solve this pair by substitution:
From 2.) we know that x = 10 - 2y
We can then substitute 10 - 2y for x in equation (1.:
5*(10-2y) + 3*y = 1
Solve for y.
Once we know what y is we can substitute that value for y in equation (1. and solve for x.